### v2

# A note on the oriented line graph operation

Abstract
Oriented line graph, introduced by Kotani and Sunada (2000), is closely related to Hashimato’s non-backtracking matrix (1989). It is known that for regular graphs $G$, the eigenvalues of the adjacency matrix of the oriented line graph $\vec{L}(G)$ of $G$ are the reciprocals of the poles of the Ihara zeta function of $G$. We determine the characteristic polynomials of the adjacency matrix of the underlying undirected graph of $\vec{L}(G)$ and the skew-symmetric adjacency matrix (and Hermitian adjacency matrix) of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$.
A locally bijective (resp. injective) homomorphism from a graph $G$ to a graph $H$ is a mapping $\psi\colon V(G)\to V(H)$ such that for every vertex $v$ of $G$, the restriction of $\psi$ to the neighborhood $N_{G}(v)$ is a bijection (resp. injection) from $N_{G}(v)$ to $N_{H}(\psi(v))$ (Fiala and Kratochvíl, 2008). An out-neighborhood bijective (resp. injective) homomorphism from a directed graph $\vec{G}$ to a directed graph $\vec{H}$ is a mapping $\psi\colon V(\vec{G})\to V(\vec{H})$ such that for every vertex $v$ of $\vec{G}$, the restriction of $\psi$ to the out-neighborhood $N_{\vec{G}}^{+}(v)$ is a bijection (resp. injection) from $N_{\vec{G}}^{+}(v)$ to $N_{\vec{H}}^{+}(\psi(v))$ (Antony and Shalu, 2025). We prove that the existence of a locally bijective (resp. injective) homomorphism from a graph $G$ of minimum degree at least 3 to a graph $H$ is equivalent to the existence of an out-neighborhood bijective (resp. injective) homomorphism from $\vec{L}(G)$ to $\vec{L}(H)$. We also prove some results on the coloring variants distance-two coloring and star coloring.
1 Introduction
The notion of oriented line graphs is introduced by Kotani and Sunada [11], and they are closely related to Hashimato’s [10] non-backtracking matrix. Let $G$ be a (simple undirected) graph. The oriented line graph of $G$ is the oriented graph with vertex set $\bigcup_{uv\in E(G)}\{(u,v),(v,u)\}$ and there is an arc from a vertex $(u,v)$ to a vertex $(v,w)$ in it when $u\neq w$ [17]. We denote the oriented line graph of $G$ by $\vec{L}(G)$, and its underlying undirected graph by $L^{*}(G)$. We can rephrase the definition of oriented line graph in terms of $\vec{D}(G)$, the simple directed graph obtained from $G$ be replacing each edge $uv$ of $G$ by two arcs $(u,v)$ and $(v,u)$ (i.e., replacing each edge by a digon). The oriented line graph of $G$ is the oriented graph with the set of arcs of $\vec{D}(G)$ as its vertex set, and there is an arc in it from a vertex $(u,v)$ to a vertex $(v,w)$ when $u\neq w$ (see Fig. 1). The non-backtracking matrix of $G$ is the adjacency matrix of $\vec{L}(G)$.
(a) $G$
(b) $\vec{D}(G)$
(c) $\vec{L}(G)$
(d) $\vec{L}(G)$ drawn on top of $\vec{D}(G)$
Figure 1: An example of the oriented line graph operation.
The characteristic polynomial of the adjacency matrix of $\vec{L}(G)$ is well-known for $d$-regular graphs $G$ with $d\geq 3$. It is known that the roots of this polynomial are the reciprocals of the poles of the Ihara zeta function of $G$ [10]. We determine the characteristic polynomial of (the adjacency matrix of) $L^{*}(G)$ and the skew-symmetric adjacency matrix of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$. We prove that the existence of certain types of locally constrained homomorphisms from a graph $G$ to a graph $H$ is equivalent to the existence of certain types of locally constrained homomorphisms from $\vec{L}(G)$ to $\vec{L}(H)$. We also prove some results on the coloring variants distance-two coloring and star coloring.
All graphs we consider are simple and finite; they are also undirected unless otherwise specified. To clarify, the directed graphs we consider are also simple (i.e., no parallel arcs, although two arcs in opposite directions could be present). We denote the first projection map (i.e., $(x,y)\mapsto x$) by $\operatorname{proj_{1}}$, and the second projection map by $\operatorname{proj_{2}}$. For convenience, we assume that the vertex set of the complete graph $K_{q}$ is $\mathbb{Z}_{q}$ for all $q\in\mathbb{Z}$. Throughout this paper, $G$ denotes a graph on which an operation, such as line graph or oriented line graph operation, is applied, and we denote the number of vertices and number of edges of $G$ by $n$ and $m$, respectively.
This paper is organized as follows. Section 2 discusses our results related to the spectrum of some matrices associated with the oriented line graph operation. Section 3 discusses the results involving graph homomorphisms, orientations and colorings. Section 3 includes the subsections 3.1 giving necessary definitions, 3.2 dealing with graph homomorphisms, and 3.3 dealing with distance-two coloring and star coloring. We conclude with Section 4.
2 Spectrum of $L^{*}(G)$
In this section, we determine the characteristic polynomial of (the adjacency matrix of) $L^{*}(G)$, and the skew-symmetric adjacency matrix and Hermitian adjacency matrix of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$.
Lubetzky and Peres [15] proved that the adjacency matrix of an oriented line graph is unitarily similar to a block-diagonal matrix with $1\times 1$ and $2\times 2$ blocks.
Theorem 2.1(Lubetzky and Peres [15]).
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the non-backtracking matrix of $G$ is unitarily similar to the following block-diagonal matrix $C$,
$C=\operatorname{diag}\left(\begin{bmatrix}\theta_{1}&\alpha_{1}\\ 0&\theta^{\prime}_{1}\end{bmatrix},\ldots,\begin{bmatrix}\theta_{n}&\alpha_{n}\\ 0&\theta^{\prime}_{n}\end{bmatrix},-1,\dots,-1,1,\dots,1\right)$ (1)
where 1-1 and $1$ are repeated mnm-n times,
$|\alpha_{i}|=\begin{cases}d-2\text{ if }|\lambda_{i}|\leq 2\sqrt{d-1}\\ \sqrt{d^{2}-\lambda_{i}^{2}}\text{ otherwise}\end{cases}$
and $\{\theta_{i},\theta_{i}^{\prime}\}=\{(\lambda/2)+\sqrt{(\lambda/2)^{2}-(d-1)},(\lambda/2)-\sqrt{(\lambda/2)^{2}-(d-1)}\}$ for $1\leq i\leq n$.
Using Theorem 2.1, we determine the characteristic polynomial of $L^{*}(G)$.
Theorem 2.2.
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the characteristic polynomial of $L^{*}(G)$ is $(x^{2}-4)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$.
Proof.
Note that $\theta_{i}+\theta_{i}^{\prime}=\lambda_{i}$, and $\theta_{i}\theta_{i}^{\prime}=d-1$ for all $i$. Let $B$ denote non-backtracking matrix of $G$; that is, the adjacency matrix of $\vec{L}(G)$. By Theorem 2.1, $B$ is unitarily similar to the matrix $C$ of Equation 1. That is, $B=P^{*}CP$ for some unitary matrix $P$ (i.e., $P^{-1}=P^{*}$). Note that $B+B^{t}$ is the adjacency matrix of $L^{*}(G)$. Also, $B+B^{t}=B+B^{*}=P^{*}CP+(P^{*}CP)^{*}=P^{*}CP+P^{*}C^{*}P=P^{*}(C+C^{*})P$. That is, $B+B^{t}$ is unitarily similar to $C+C^{*}$.
$C^{*}=\operatorname{diag}\left(\begin{bmatrix}\mkern 1.5mu\overline{\mkern-1.5mu\theta_{1}\mkern-1.5mu}\mkern 1.5mu&0\\ \mkern 1.5mu\overline{\mkern-1.5mu\alpha_{1}\mkern-1.5mu}\mkern 1.5mu&\mkern 1.5mu\overline{\mkern-1.5mu\theta_{1}^{\prime}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix},\dots,\begin{bmatrix}\mkern 1.5mu\overline{\mkern-1.5mu\theta_{n}\mkern-1.5mu}\mkern 1.5mu&0\\ \mkern 1.5mu\overline{\mkern-1.5mu\alpha_{n}\mkern-1.5mu}\mkern 1.5mu&\mkern 1.5mu\overline{\mkern-1.5mu\theta_{n}^{\prime}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix},-1,\dots,-1,1,\dots,1\right)\text{, and }$
$C+C^{*}=\operatorname{diag}\left(\begin{bmatrix}\theta_{1}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{1}\mkern-1.5mu}\mkern 1.5mu&\alpha_{1}\\ \mkern 1.5mu\overline{\mkern-1.5mu\alpha_{1}\mkern-1.5mu}\mkern 1.5mu&\theta_{1}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{1}^{\prime}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix},\dots,\begin{bmatrix}\theta_{n}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{n}\mkern-1.5mu}\mkern 1.5mu&\alpha_{n}\\ \mkern 1.5mu\overline{\mkern-1.5mu\alpha_{n}\mkern-1.5mu}\mkern 1.5mu&\theta_{n}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{n}^{\prime}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix},-2,\dots,-2,2,\dots,2\right)$,
where 1-1 and $1$ (resp. 2-2 and $2$) are repeated mnm-n times. If $|\lambda_{i}|\leq 2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=\lambda_{i}=\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu$. If $|\lambda_{i}|>2\sqrt{d-1}$, then $\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}$, and $\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu=2\theta_{i}^{\prime}$.
Since $B+B^{t}$ is unitarily similar to $C+C^{*}$, the required polynomial $\operatorname{char}(B+B^{t};x)=\operatorname{char}(C+C^{*};x)=det(xI-(C+C^{*}))=(x+2)^{m-n}(x-2)^{m-n}\prod_{i=1}^{n}h_{i}(x)$,
where $h_{i}(x)=\left((x-(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu))(x-(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu))-\alpha_{i}\mkern 1.5mu\overline{\mkern-1.5mu\alpha_{i}\mkern-1.5mu}\mkern 1.5mu\right)$. Note that $h_{i}(x)=x^{2}-x(\theta_{i}+\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}+\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)+(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu)(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)-|\alpha_{i}|^{2}=x^{2}-x(\lambda_{i}+\lambda_{i})+(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu)(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)-|\alpha_{i}|^{2}$ That is, $h_{i}(x)=x^{2}-2\lambda_{i}x+(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu)(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)-|\alpha_{i}|^{2}$. Clearly, $(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu)(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)=\lambda_{i}^{2}$ if $|\lambda_{i}|\leq 2\sqrt{d-1}$, and $(\theta_{i}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}\mkern-1.5mu}\mkern 1.5mu)(\theta_{i}^{\prime}+\mkern 1.5mu\overline{\mkern-1.5mu\theta_{i}^{\prime}\mkern-1.5mu}\mkern 1.5mu)=4\theta_{i}\theta_{i}^{\prime}$ otherwise. We know that $|\alpha_{i}|^{2}=(d-2)^{2}$ if $|\lambda_{i}|\leq 2\sqrt{d-1}$ and $|\alpha_{i}|^{2}=d^{2}-\lambda_{i}^{2}$ otherwise. Therefore,
$h_{i}(x)=\begin{cases}x^{2}-2\lambda_{i}x+\lambda_{i}^{2}-(d-2)^{2}\text{ if }|\lambda_{i}|\leq 2\sqrt{d-1}\\ x^{2}-2\lambda_{i}x+4(d-1)-(d^{2}-\lambda_{i}^{2})\text{ otherwise}\end{cases}$
Thus, in both cases, $h_{i}(x)=x^{2}-2\lambda_{i}x+\lambda_{i}^{2}-(d-2)^{2}=(x-\lambda_{i})^{2}-(d-2)^{2}$. This completes the proof. ∎
Remark: $\operatorname{char}(L(G);x)=(x+2)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})-(d-2)\right)$, and
$\operatorname{char}(L(\vec{D}(G));x)=x^{m-n}\prod_{i=1}^{n}(x-\lambda_{i})$ [4] (note that $L(\vec{H})$ denotes the line digraph of a digraph $\vec{H}$).
Corollary 1.
If a connected $d$-regular graph $G$ is integral and $d\geq 3$, then $L^{*}(G)$ is integral.
Note: A graph is said to be integral if all of its eigenvalues are integers.
The skew-symmetric adjacency matrix of $\vec{L}(G)$ can be determined similarly. The proof is omitted (hint: $B-B^{t}=P^{*}(C-C^{*})P$).
Theorem 2.3.
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Then, the characteristic polynomial of the skew-symmetric adjacency matrix of $\vec{L}(G)$ is $x^{2(m-n)}\prod_{i=1}^{n}(x^{2}+d^{2}-\lambda_{i}^{2})$. ∎
Remark: Therefore, the characteristic polynomial of the Hermitian adjacency matrix of $\vec{L}(G)$ is $x^{2(m-n)}\prod_{i=1}^{n}(x^{2}+\lambda_{i}^{2}-d^{2})$.
The next theorem follows from Lemma 3.1 of Bilu and Linial [5] since $L^{*}(G)$ is a 2-lift of $L(G)$ [3], and the characteristic polynomials of $L^{*}(G)$ and $L(G)$ are
$(x^{2}-4)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})^{2}-(d-2)^{2}\right)$ and $(x+2)^{m-n}\prod_{i=1}^{n}\left((x-\lambda_{i})-(d-2)\right)$,
respectively.
Theorem 2.4.
Let $G$ be a connected $d$-regular graph with $n$ vertices, $m$ edges and eigenvalues $\lambda_{1},\lambda_{2},\dots,\lambda_{n}$, where $d\geq 3$. Let $\psi$ be an LBH from $L^{*}(G)$ to $L(G)$. Let $\{V_{0},V_{1}\}$ be a partition of the vertex set of $\vec{L}(G)$ such that $|\psi^{-1}(v)\cap V_{i}|=1$ for each vertex $v$ of $G$ and each $i\in\mathbb{Z}_{2}$. Let $(L(G),s)$ be the signed graph such that for each edge $uv$ of $L(G)$, we have $s(uv)=1$ if and only if $uv$ belongs to the cut $(V_{0},V_{1})$ in $\vec{L}(G)$. Then, the characteristic polynomial of the signed adjacency matrix of $(L(G),s)$ is $(x-2)^{m-n}\prod_{i=1}^{n}(x-\lambda_{i}+d-2)$. ∎
Remark: The existence and uniqueness of $(L(G),s)$ is guaranteed (uniqueness up to switching).
3 Colorings, Orientations and Homomorphisms
In this section, we prove that the existence of certain types of homomorphisms from a graph $G$ to a graph $H$ is equivalent to the existence of certain types of homomorphisms from $\vec{L}(G)$ to $\vec{L}(H)$. We also prove some results on the coloring variants distance-two coloring and star coloring defined below.
3.1 Definitions
Let $G$ be a graph, and let $q\in\mathbb{Z}^{+}$. A (proper) coloring of $G$ is a function from $V(G)$ to a set $S$ such that $f(u)\neq f(v)$ for every edge $uv$ of $G$ (the convention is to take a subset of $\mathbb{Z}$ as $S$, and $S$ is often of the form $\{1,2,\dots,k\}$ for some $k\in\mathbb{Z}^{+}$). A $q$-coloring of $G$ is a coloring $f$ of $G$ such that $|f(V(G))|\leq q$. Throughout this paper, we take $\mathbb{Z}_{q}$ as the co-domain of a $q$-coloring $f$ of $G$ (i.e., $f\colon V(G)\to\mathbb{Z}_{q}$).
A $q$-colored graph is an ordered pair $(G,f)$ such that $G$ is a graph, and $f$ is a $q$-coloring of $G$. A distance-two $q$-coloring of $G$ is a $q$-coloring $f$ of $G$ such that there is no bicolored 3-vertex path in $(G,f)$ [12]. A star $q$-coloring of $G$ is a $q$-coloring $f$ of $G$ such that there is no bicolored 4-vertex path in $(G,f)$ [9].
Next, we define two special cases of colored orientations. A $q$-colored simple directed graph $(\vec{G},f)$ is a $q$-colored in-orientation (of a multigraph) if for every vertex $v$ of $(\vec{G},f)$, (i) no out-neighbor of $v$ has the same color as an in-neighbor of $v$, and (ii) out-neighbors of $v$ have pairwise distinct colors [6]. A $q$-colored in-orientation is a $q$-colored Monochromatic In-neighborhood In-orientation (in short, $q$-colored MINI-orientation) if for every vertex $v$ of $(\vec{G},f)$, all in-neighbors of $v$ have the same color [3]. A coloring $f$ of a graph $G$ is a star coloring of $G$ if and only if there exists an orientation $\vec{G}$ of $G$ such that $(\vec{G},f)$ is a colored in-orientation [1, 16]. The notion of $q$-colored MINI-orientations was introduced in [3] to study star coloring of regular graphs.
A locally bijective (resp. injective) homomorphism from a graph $G$ to a graph $H$ is a mapping $\psi\colon V(G)\to V(H)$ such that for every vertex $v$ of $G$, the restriction of $\psi$ to the neighborhood $N_{G}(v)$ is a bijection (resp. injection) from $N_{G}(v)$ to $N_{H}(\psi(v))$ [8]. A locally bijective (resp. injective) homomorphism from a directed graph $\vec{G}$ to a directed graph $\vec{H}$ is a mapping $\psi\colon V(\vec{G})\to V(\vec{H})$ such that for every vertex $v$ of $\vec{G}$, (i) the restriction of $\psi$ to the out-neighborhood $N_{\vec{G}}^{+}(v)$ is a bijection (resp. injection) from $N_{\vec{G}}^{+}(v)$ to $N_{\vec{H}}^{+}(\psi(v))$, and (ii) the restriction of $\psi$ to the in-neighborhood $N_{\vec{G}}^{-}(v)$ is a bijection (resp. injection) from $N_{\vec{G}}^{-}(v)$ to $N_{\vec{H}}^{-}(\psi(v))$ (this definition of locally bijective homomorphism between directed graphs is well-known, whereas locally injective homomorphisms between directed graphs appears to be new). An out-neighborhood bijective (resp. injective) homomorphism from a directed graph $\vec{G}$ to a directed graph $\vec{H}$ is a mapping $\psi\colon V(\vec{G})\to V(\vec{H})$ such that for every vertex $v$ of $\vec{G}$, the restriction of $\psi$ to the out-neighborhood $N_{\vec{G}}^{+}(v)$ is a bijection (resp. injection) from $N_{\vec{G}}^{+}(v)$ to $N_{\vec{H}}^{+}(\psi(v))$ [2, 3]. We abbreviate Locally Bijective Homomorphism as LBH, Locally Injective Homomorphism as LIH, Out-neighborhood Bijective Homomorphism as OBH, and Out-neighborhood Injective Homomorphism as OIH.
3.2 Locally constrained homomorphisms
In this subsection, we present two results that link Locally Bijective/Injective Homomorphism (LBH/LIH) to Out-neighborhood Bijective/Injective Homomorphism (OBH/OIH) by means of the oriented line graph operation.
Theorem 3.1.
Let $G$ and $H$ be two connected graphs with minimum degree at least 3. Then, there is a bijection from the set of LBHs from $G$ to $H$ onto the set of OBHs from $\vec{L}(G)$ to $\vec{L}(H)$. Moreover, every OBH from $\vec{L}(G)$ to $\vec{L}(H)$ is an LBH from $\vec{L}(G)$ to $\vec{L}(H)$.
Proof.
First, we define a mapping $\varphi$ from the set of LBHs from $G$ to $H$ to the set of OBHs from $\vec{L}(G)$ to $\vec{L}(H)$. For each LBH $\psi$ from $G$ to $H$, its image $\varphi(\psi)$ is the mapping $\psi^{\prime}\colon V(\vec{L}(G))\to V(\vec{L}(H))$ defined as $\psi^{\prime}((u,v))=(\psi(u),\psi(v))$ for each $(u,v)\in V(\vec{L}(G))$. The mapping $\psi^{\prime}$ is well-defined since $\psi$ is a homomorphism from $G$ to $H$. Let $\psi$ be an LBH from $G$ to $H$, and let $\psi^{\prime}=\varphi(\psi)$. To prove that $\psi^{\prime}$ is an OBH from $\vec{L}(G)$ to $\vec{L}(H)$, consider an arbitrary vertex $(u,v)$ of $\vec{L}(G)$ and the out-neighbors of $(u,v)$ in $\vec{L}(G)$, say $(v,x_{1}),(v,x_{2}),\dots,(v,x_{q})$, where $q\geq 2$ (because $\delta(G)\geq 3$). Then, the neighbors of $v$ in $G$ are $u,x_{1},\dots,x_{q}$. Since $\psi$ is an LBH from $G$ to $H$, the neighbors of $\psi(v)$ in $H$ are $\psi(u),\psi(x_{1}),\dots,\psi(x_{q})$ (and they are pairwise distinct). Hence, the out-neighbors of $(\psi(u),\psi(v))$ in $\vec{L}(H)$ are $(\psi(v),\psi(x_{1})),\dots,(\psi(v),\psi(x_{q}))$. Thus, the restriction of $\psi^{\prime}$ to the out-neighborhood of $(u,v)$ is a bijection from $N^{+}_{\vec{L}(G)}(u,v)$ onto $N^{+}_{\vec{L}(H)}(\psi^{\prime}((u,v)))$. Since $(u,v)$ is arbitrary, $\psi^{\prime}$ is indeed an OBH from $\vec{L}(G)$ to $\vec{L}(H)$. In fact, $\psi^{\prime}$ is an LBH from $\vec{L}(G)$ to $\vec{L}(H)$; the proof for in-neighborhood is similar.
Claim 1:
For each LBH $\psi$ from $G$ to $H$, the mapping $\varphi(\psi)$ is an LBH from $\vec{L}(G)$ to $\vec{L}(H)$.
This proves that (i) $\varphi$ is a mapping from the set of LBHs from $G$ to $H$ to the set of OBHs from $\vec{L}(G)$ to $\vec{L}(H)$, and (ii) the range of $\varphi$ is a subset of the set of LBHs from $\vec{L}(G)$ to $\vec{L}(H)$.
Next, we exhibit a mapping $\varphi^{\prime}$ from the set of OBHs from $\vec{L}(G)$ to $\vec{L}(H)$ to the set of LBHs from $G$ to $H$.
Claim 2:
For each OBH $\psi^{\prime}$ from $\vec{L}(G)$ to $\vec{L}(H)$ and for each pair of vertices $(u,v)$ and $(w,v)$ of $\vec{L}(G)$, we have $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$.
Suppose that $\psi^{\prime}$ is an OBH from $\vec{L}(G)$ to $\vec{L}(H)$. Let $(u,v)$ and $(w,v)$ be vertices of $\vec{L}(G)$. Then, $u$ and $w$ are (distinct) neighbors of $v$ in $G$. Since $\delta(G)\geq 3$, there is a neighbor $x$ of $v$ in $G$ different from $u$ and $w$. Clearly, $(v,x)$ in an out-neighbor of both $(u,v)$ and $(w,v)$ in $\vec{L}(G)$. Since $\psi^{\prime}$ is an OBH from $\vec{L}(G)$ to $\vec{L}(H)$, the vertex $\psi^{\prime}((v,x))$ is an out-neighbor of both $\psi^{\prime}((u,v))$ and $\psi^{\prime}((w,v))$ in $\vec{L}(H)$. Thus, by the definition of oriented line graph, $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{1}}(\psi^{\prime}((v,x)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$. Since $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$, Claim 2 is proved.
For each OBH $\psi^{\prime}$ from $\vec{L}(G)$ to $\vec{L}(H)$, its image $\varphi^{\prime}(\psi^{\prime})$ is a mapping $\psi\colon V(G)\to V(H)$ defined as $\psi(v)=\operatorname{proj_{2}}(\psi^{\prime}((u,v)))$ for some $(u,v)\in V(\vec{L}(G))$. By Claim 2, $\psi$ is well-defined. Let $\psi^{\prime}$ be an OBH from $\vec{L}(G)$ to $\vec{L}(H)$, and let $\psi=\varphi^{\prime}(\psi^{\prime})$. To prove that $\psi$ is an LBH from $G$ to $H$, consider an arbitrary vertex $v$ of $G$, and its neighbors in $G$, say $u_{1},u_{2},\dots,u_{t}$, where $t\geq 3$. Clearly, for $1\leq i\leq t$, the set of out-neighbors of $(u_{i},v)$ in $\vec{L}(G)$ is $\{(v,u_{j})\colon 1\leq j\leq t\text{ and }j\neq i\}$.
Since $\psi^{\prime}$ is a homomorphism from $\vec{L}(G)$ to $\vec{L}(H)$, the vertex $\psi^{\prime}((v,u_{2}))$ is an out-neighbor of $\psi^{\prime}((u_{1},v))$. Hence, by definitions of oriented line graph and $\psi$, apparently $(\operatorname{proj_{1}}\circ\psi^{\prime})(v,u_{2})=(\operatorname{proj_{2}}\circ\psi^{\prime})(u_{1},v)=\psi(v)$. Hence, $\psi^{\prime}((v,u_{2}))=(\psi(v),\psi(u_{2}))$. Similarly, $\psi^{\prime}((v,u_{j}))=(\psi(v),\psi(u_{j}))$ for $1\leq j\leq t$. Since the restriction of $\psi^{\prime}$ to $N^{+}_{\vec{L}(G)}((u_{1},v))$ is a bijection from $N^{+}_{\vec{L}(G)}((u_{1},v))$ to the set $N^{+}_{\vec{L}(H)}(\psi^{\prime}((u_{1},v)))$, the out-neighbors of $\psi^{\prime}((u_{1},v))$ in $\vec{L}(H)$ are $\psi^{\prime}((v,u_{2})),\dots,\allowbreak\psi^{\prime}((v,u_{t}))$, and they are pairwise distinct. That is, the out-neighbors of $\psi^{\prime}((u_{1},v))$ in $\vec{L}(H)$ are $(\psi(v),\psi(u_{2})),\dots,\allowbreak(\psi(v),\psi(u_{t}))$, and they are pairwise distinct. Similarly, the out-neighbors of $\psi^{\prime}((u_{t},v))$ in $\vec{L}(H)$ are $\psi^{\prime}((v,u_{1})),\dots,\psi^{\prime}((v,u_{t-1}))$, and they are pairwise distinct. Hence, by the definition of oriented line graph, the neighbors of $\psi(v)$ in $H$ are $\psi(u_{1}),\psi(u_{2}),\dots,\psi(u_{t})$. Since $(\psi(v),\psi(u_{2})),\dots,\allowbreak(\psi(v),\psi(u_{t}))$ are pairwise distinct, $\psi(u_{2}),\dots,\psi(u_{t})$ are pairwise distinct. Similarly, $\psi(u_{1}),\dots,\psi(u_{t-1})$ are pairwise distinct (consider $\psi^{\prime}$ restricted to the set $N^{+}_{\vec{L}(G)}(u_{t},v)$), and $\psi(u_{1}),\psi(u_{3}),\psi(u_{4}),\dots,\psi(u_{t})$ are pairwise distinct (consider $\psi^{\prime}$ restricted to $N^{+}_{\vec{L}(G)}(u_{2},v)$). Therefore, $\psi(u_{1}),\psi(u_{2}),\dots,\psi(u_{t})$, the neighbors of $\psi(v)$ in $H$, are pairwise distinct. This proves that the restriction of $\psi$ to $N_{G}(v)$ is a injection from $N_{G}(v)$ to $N_{H}(\psi(v))$. Since $v$ is arbitrary, $\psi$ is an LIH from $G$ to $H$. This proves that $\varphi^{\prime}$ is a mapping from the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$ to the set of LIHs from $G$ to $H$.
It is easy to verify that $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are identity maps. Therefore, $\varphi$ and $\varphi^{\prime}$ are bijections. Since $\varphi(\psi)$ is an LIH from $\vec{L}(G)$ to $\vec{L}(H)$ for every LIH $\psi$ from $G$ to $H$ (see Claim 1), the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$ is equal to the set of LIHs from $\vec{L}(G)$ to $\vec{L}(H)$. ∎
Theorem 3.2.
Let $G$ and $H$ be two connected graphs with minimum degree at least 3. Then, there is a bijection from the set of LIHs from $G$ to $H$ onto the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$. Moreover, every OIH from $\vec{L}(G)$ to $\vec{L}(H)$ is an LIH from $\vec{L}(G)$ to $\vec{L}(H)$.
Proof.
First, we define a mapping $\varphi$ from the set of LIHs from $G$ to $H$ to the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$. For each LIH $\psi$ from $G$ to $H$, its image $\varphi(\psi)$ is the mapping $\psi^{\prime}\colon V(\vec{L}(G))\to V(\vec{L}(H))$ defined as $\psi^{\prime}((u,v))=(\psi(u),\psi(v))$ for each $(u,v)\in V(\vec{L}(G))$. The mapping $\psi^{\prime}$ is well-defined since $\psi$ is a homomorphism from $G$ to $H$. Let $\psi$ be an LIH from $G$ to $H$. To prove that $\psi^{\prime}$ is an OIH from $\vec{L}(G)$ to $\vec{L}(H)$, consider an arbitrary vertex $(u,v)$ of $\vec{L}(G)$. Let $(v,x_{1}),(v,x_{2}),\dots,(v,x_{q})$ be the out-neighbors of $(u,v)$ in $\vec{L}(G)$, where $q\geq 0$. Then, the neighbors of $v$ in $G$ are $u,x_{1},\dots,x_{q}$, and $q\geq 2$ (because $\delta(G)\geq 3$). Since $\psi$ is an LIH from $G$ to $H$, the neighbors of $\psi(v)$ in $H$ include $\psi(u),\psi(x_{1}),\dots,\psi(x_{q})$, and they are pairwise distinct. Hence, the out-neighbors of $(\psi(u),\psi(v))$ in $\vec{L}(H)$ include $(\psi(v),\psi(x_{1})),\dots,(\psi(v),\psi(x_{q}))$, and they are pairwise distinct. Thus, the restriction of $\psi^{\prime}$ to the out-neighborhood of $(u,v)$ is an injection from $N^{+}_{\vec{L}(G)}(u,v)$ onto $N^{+}_{\vec{L}(H)}(\psi^{\prime}((u,v)))$. Since $(u,v)$ is arbitrary, $\psi^{\prime}$ is indeed an OIH from $\vec{L}(G)$ to $\vec{L}(H)$. In fact, $\psi^{\prime}$ is an LIH from $\vec{L}(G)$ to $\vec{L}(H)$; the proof for in-neighborhood is similar.
Claim 3:
For each LIH $\psi$ from $G$ to $H$, the mapping $\varphi(\psi)$ is an LIH from $\vec{L}(G)$ to $\vec{L}(H)$.
This proves that (i) $\varphi$ is a mapping from the set of LIHs from $G$ to $H$ to the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$, and (ii) the range of $\varphi$ is a subset of the set of LIHs from $\vec{L}(G)$ to $\vec{L}(H)$.
Next, we exhibit a mapping $\varphi^{\prime}$ from the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$ to the set of LIHs from $G$ to $H$.
Claim 4:
For each OIH $\psi^{\prime}$ from $\vec{L}(G)$ to $\vec{L}(H)$ and for each pair of vertices $(u,v)$ and $(w,v)$ of $\vec{L}(G)$, we have $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$.
Suppose that $\psi^{\prime}$ is an OIH from $\vec{L}(G)$ to $\vec{L}(H)$. Let $(u,v)$ and $(w,v)$ be vertices of $\vec{L}(G)$. Then, $u$ and $w$ are (distinct) neighbors of $v$ in $G$. Since $\delta(G)\geq 3$, there is a neighbor $x$ of $v$ in $G$ different from $u$ and $w$. Clearly, $(v,x)$ in an out-neighbor of both $(u,v)$ and $(w,v)$ in $\vec{L}(G)$. Since $\psi^{\prime}$ is an OIH from $\vec{L}(G)$ to $\vec{L}(H)$, the vertex $\psi^{\prime}((v,x))$ is an out-neighbor of both $\psi^{\prime}((u,v))$ and $\psi^{\prime}((w,v))$ in $\vec{L}(H)$. Thus, by the definition of oriented line graph, $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{1}}(\psi^{\prime}((v,x)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$. Since $\operatorname{proj_{2}}(\psi^{\prime}((u,v)))=\operatorname{proj_{2}}(\psi^{\prime}((w,v)))$, Claim 4 is proved.
For each OIH $\psi^{\prime}$ from $\vec{L}(G)$ to $\vec{L}(H)$, its image $\varphi^{\prime}(\psi^{\prime})$ is a mapping $\psi\colon V(G)\to V(H)$ defined as $\psi(v)=\operatorname{proj_{2}}(\psi^{\prime}((u,v)))$ for some $(u,v)\in V(\vec{L}(G))$. By Claim 4, $\psi$ is well-defined. Let $\psi^{\prime}$ be an OIH from $\vec{L}(G)$ to $\vec{L}(H)$, and let $\psi=\varphi^{\prime}(\psi^{\prime})$. To prove that $\psi$ is an LBH from $G$ to $H$, consider an arbitrary vertex $v$ of $G$, and its neighbors in $G$, say $u_{1},u_{2},\dots,u_{t}$, where $t\geq 3$. Clearly, for $1\leq i\leq t$, the set of out-neighbors of $(u_{i},v)$ in $\vec{L}(G)$ is $\{(v,u_{j})\colon 1\leq j\leq t\text{ and }j\neq i\}$.
Since $\psi^{\prime}$ is a homomorphism from $\vec{L}(G)$ to $\vec{L}(H)$, the vertex $\psi^{\prime}((v,u_{2}))$ is an out-neighbor of $\psi^{\prime}((u_{1},v))$. Hence, by definitions of oriented line graph and $\psi$, apparently $(\operatorname{proj_{1}}\circ\psi^{\prime})(v,u_{2})=(\operatorname{proj_{2}}\circ\psi^{\prime})(u_{1},v)=\psi(v)$. Hence, $\psi^{\prime}((v,u_{2}))=(\psi(v),\psi(u_{2}))$. Similarly, $\psi^{\prime}((v,u_{j}))=(\psi(v),\psi(u_{j}))$ for $1\leq j\leq t$. Since the restriction of $\psi^{\prime}$ to $N^{+}_{\vec{L}(G)}((u_{1},v))$ is a injection from $N^{+}_{\vec{L}(G)}((u_{1},v))$ to the set $N^{+}_{\vec{L}(H)}(\psi^{\prime}((u_{1},v)))$, the out-neighbors of $\psi^{\prime}((u_{1},v))$ in $\vec{L}(H)$ include $\psi^{\prime}((v,u_{2})),\allowbreak\dots,\psi^{\prime}((v,u_{t}))$, and they are pairwise distinct. That is, the out-neighbors of $\psi^{\prime}((u_{1},v))$ in $\vec{L}(H)$ include $(\psi(v),\psi(u_{2})),\dots,(\psi(v),\psi(u_{t}))$, and they are pairwise distinct. Similarly, the out-neighbors of $\psi^{\prime}((u_{t},v))$ in $\vec{L}(H)$ include vertices $\psi^{\prime}((v,u_{1})),\allowbreak\dots,\psi^{\prime}((v,u_{t-1}))$, and they are pairwise distinct. Hence, by definition of oriented line graph, the neighbors of $\psi(v)$ in $H$ include $\psi(u_{1}),\psi(u_{2}),\dots,\psi(u_{t})$. Since $(\psi(v),\psi(u_{2})),\dots,\allowbreak(\psi(v),\psi(u_{t}))$ are pairwise distinct, $\psi(u_{2}),\dots,\psi(u_{t})$ are pairwise distinct. Similarly, $\psi(u_{1}),\dots,\psi(u_{t-1})$ are pairwise distinct (consider $\psi^{\prime}$ restricted to $N^{+}_{\vec{L}(G)}(u_{t},v)$), and $\psi(u_{1}),\psi(u_{3}),\psi(u_{4}),\dots,\psi(u_{t})$ are pairwise distinct (consider $\psi^{\prime}$ restricted to $N^{+}_{\vec{L}(G)}(u_{2},v)$). Therefore, the neighbors of $\psi(v)$ in $H$ include $\psi(u_{1}),\psi(u_{2}),\dots,\psi(u_{t})$, and they are pairwise distinct. This proves that the restriction of $\psi$ to $N_{G}(v)$ is a bijection from $N_{G}(v)$ to $N_{H}(\psi(v))$. Since $v$ is arbitrary, $\psi$ is an LBH from $G$ to $H$. This proves that $\varphi^{\prime}$ is a mapping from the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$ to the set of LBHs from $G$ to $H$.
It is easy to verify that $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are identity maps. Therefore, $\varphi$ and $\varphi^{\prime}$ are bijections. Since $\varphi(\psi)$ is an LIH from $\vec{L}(G)$ to $\vec{L}(H)$ for every LIH $\psi$ from $G$ to $H$ (see Claim 3), the set of OIHs from $\vec{L}(G)$ to $\vec{L}(H)$ is equal to the set of LIHs from $\vec{L}(G)$ to $\vec{L}(H)$. ∎
3.3 Distance-two coloring and star coloring
In this subsection, we characterize distance $q$-colorable graphs $G$ with minimum degree at least 3 in terms of colored MINI-orientations of $\vec{L}(G)$.
Theorem 3.3.
Let $G$ be graph with the minimum degree $\delta(G)\geq 3$, and let $q\in\mathbb{N}$. Then, the set of distance-two $q$-colorings of $G$ and the set of $q$-colored MINI-orientations of $\vec{L}(G)$ have the same cardinality.
Proof.
First, we define a mapping $\varphi$ from the set of distance-two $q$-colorings of $G$ to the set of $q$-colored MINI-orientations of $\vec{L}(G)$ as $\psi(f)=(\vec{L}(G),f\circ\mathrm{proj}_{2})$. Suppose that $f$ be a distance-two $q$-coloring of $G$. Define $h\colon V(\vec{L}(G))\to\mathbb{Z}_{q}$ as $h=f\circ\mathrm{proj}_{2}$. Let $(u,v)$ be an arbitrary vertex of $\vec{L}(G)$. Let $(w_{1},u),(w_{2},u),\dots,\allowbreak(w_{s},u)$ be the in-neighbors of $(u,v)$ and $(v,x_{1}),(v,x_{2}),\dots,(v,x_{t})$ be the out-neighbors of $(u,v)$ in $\vec{L}(G)$, where $s,t\in\mathbb{N}$. By the definitions of $h$ and $\vec{L}(G)$, we have $h((u,v))=f(v)$, $h((w_{1},u))=\dots=h((w_{s},u))=f(u)$ and $h((v,x_{j}))=f(x_{j})$ for $1\leq j\leq t$. Note that $u,v,x_{1},x_{2},\dots,x_{t}$ have pairwise distinct colors under $f$. As a result, $(u,v)$ and its out-neighbors have pairwise distinct colors under $h$. Moreover, all in-neighbors of $(u,v)$ have the same color under $h$, and that color differs from that of $(u,v)$ and its out-neighbors. Since $(u,v)$ is arbitrary, $(\vec{L}(G),f\circ\mathrm{proj}_{2})$ is a $q$-colored MINI-orientation.
Next, we exhibit a mapping $\varphi^{\prime}$ from the set of $q$-colored MINI-orientations of $\vec{L}(G)$ to the set of distance-two $q$-colorings of $G$. Suppose that $(\vec{L}(G),h)$ is a $q$-colored MINI-orientation.
Claim 5:
For each vertex $v$ of $G$, there exists a unique color $c_{h,v}\in\mathbb{Z}_{q}$ such that $h((u,v))=c_{h,v}$ for every vertex of $\vec{L}(G)$ with $v$ as its second coordinate.
Let $v$ be a vertex of $G$, and let $u_{1},u_{2},\dots,u_{t}$ be the neighbors of $v$ in $G$, where $t\geq 3$. The vertices of $\vec{L}(G)$ with $v$ as its second coordinate are $(u_{1},v),(u_{2},v),\dots(u_{t},v)$. Since $(u_{2},v),\dots,(u_{t},v)$ are in-neighbors of $(v,u_{1})$, we have $h((u_{2},v))=\dots=h((u_{t},v))$. Since $(u_{1},v),\dots,(u_{t-1},v)$ are in-neighbors of $(v,u_{t})$, we have $h((u_{1},v))=\dots=h((u_{t-1},v))$. Since $t\geq 3$, there exists a unique color $c_{h,v}$ such that $h((u_{i},v))=c_{h,v}$ for $1\leq i\leq t$. This proves Claim 5.
Define $f_{h}\colon V(G)\to\mathbb{Z}_{q}$ as $f_{h}(v)=c_{h,v}$. It follows from the definitions of $c_{h,v}$ and $f_{h}$ that $f_{h}\circ\mathrm{proj}_{2}=h$. To prove that $f_{h}$ is a distance-two coloring of $G$, consider an arbitrary vertex $v$ of $G$ and two neighbors $u$ and $w$ of $v$ in $G$. Note that $f_{h}(u)=h((v,u))$, $f_{h}(v)=h((u,v))=h((w,v))$ and $f_{h}(w)=h((v,w))$. Since $h$ is a coloring of $L^{*}(G)$, we have $h((u,v))\neq h((v,w))$ and $h((v,u))\neq h((w,v))$. That is, $f_{h}(v)\neq f_{h}(w)$ and $f_{h}(u)\neq f_{h}(v)$. This show that $f_{h}$ is a coloring of $G$. Since $\delta(G)\geq 3$, the vertex $v$ of $G$ has a neighbor $x$ different from $u$ and $w$. Since $(v,u)$ and $(v,w)$ are out-neighbors of $(x,v)$ and $(\vec{L}(G),h)$ is a colored MINI-orientation, $h((v,u))\neq h((v,w))$. That is, $f_{h}(u)\neq f_{h}(w)$. This proves that $f_{h}$ is a distance-two $q$-coloring of $G$. Therefore, a mapping $\varphi^{\prime}$ defined as $(\vec{L}(G),h)\mapsto f_{h}$ is indeed a mapping from the set of $q$-colored MINI-orientations of $\vec{L}(G)$ to the set of distance-two $q$-colorings of $G$.
It is easy to verify that $\varphi\circ\varphi^{\prime}$ and $\varphi^{\prime}\circ\varphi$ are identity maps. Therefore, $\varphi$ and $\varphi^{\prime}$ are bijections. ∎
Corollary 2.
For each integer $q\geq 4$, a graph $G$ of minimum degree at least 3 is distance-two $q$-colorable if and only if $\vec{L}(G)$ admits a $q$-colored MINI-orientation. ∎
A graph admitting a $q$-colored MINI-orientation implies the graph admitting a $q$-colored in-orientation, and thus a star $q$-coloring. Thus, we have the following corollary.
Corollary 3.
$\chi_{s}(L^{*}(G))\leq\chi(G^{2})$ for every graph $G$ with minimum degree at least 3. ∎
Remark: For a graph $H$, the star chromatic number $\chi_{s}(H)$ (resp. distance-two chromatic number $\chi(H^{2})$) is the least integer $k$ such that $H$ is star $k$-colorable (resp. distance-two $k$-colorable).
Let us focus on star coloring now. For $p\geq 2$, a $2p$-regular graph is star $(p+2)$-colorable if and only if $G$ admits an out-neighborhood bijective homomorphism to $L^{*}(K_{p+2})$ [3]. Interestingly, star coloring of general graphs can be characterized directly in terms of constrained graph homomorphisms with the help of the following lemma.
Lemma 1.
A $q$-colored orientation $(\vec{G},f)$ of a graph $G$ is a $q$-colored in-orienta-
tion if and only if $f$ is an out-neighborhood injective homomorphism from $\vec{G}$ to $\vec{D}(K_{q})$. ∎
Theorem 3.4.
For $q\in\mathbb{N}$, a star $q$-coloring of a graph $G$ is an out-neighborhood injective homomorphism from an orientation of $G$ to $\vec{D}(K_{q})$. ∎
By Theorem 2.2, the characteristic polynomial of $L^{*}(K_{q})$ is $(x-2)^{(q-2)(q-1)/2}\allowbreak(x+2)^{(q-3)q/2}(x-2q+4)(x-q+4)^{q-1}(x+q-2)^{q-1}$. For $q=p+2$, the polynomial is $(x-2)^{p(p+1)/2}(x+2)^{(p-1)(p+2)/2}(x-2p)(x-p+2)^{p+1}(x+p)^{p+1}$. In [3], it is proved that for $p\geq 2$, a $K_{1,p+1}$-free $2p$-regular graph $G$ is star $(p+2)$-colorable if and only if $G$ admits an LBH to $L^{*}(K_{p+2})$. Since existence of LBH from a graph $H$ to a graph $J$ implies that the characteristic polynomial of $J$ divides the characteristic polynomial of $H$ [8], we have the following theorem.
Theorem 3.5.
Let $G$ be a $K_{1,p+1}$-free $2p$-regular star $(p+2)$-colorable graph, where $p\geq 2$. Then, the characteristic polynomial of ((the adjacency matrix of)) $G$ is divisible by $(x-2)^{p(p+1)/2}(x+2)^{(p-1)(p+2)/2}(x-p+2)^{p+1}(x+p)^{p+1}$. ∎
Consider a fixed integer $p\geq 2$. It is NP-complete to check whether a $(p+1)$-regular graph $G$ admits an LBH to $K_{p+2}$ [13] (i.e., whether $G$ is distance-two $(p+2)$-colorable [7, 14]). Hence, we have the following due to Corollary 2.
Theorem 3.6.
For $p\geq 2$, given a $(p+1)$-regular graph $G$, it is NP-complete to check whether $\vec{L}(G)$ admits a $(p+2)$-colored MINI-orientation.
Thus, we have the following lemma.
Lemma 2.
For each $p\geq 2$, given an orientation $\vec{G}$ of a $2p$-regular graph, it is NP-complete to check whether $\vec{G}$ admits a $(p+2)$-colored MINI-orientation. ∎
Since every $(p+2)$-colored in-orientation of a $2p$-regular graph (with $p\geq 2$) is a $(p+2)$-colored MINI-orientation [3], we can rephrase the lemma as follows.
Theorem 3.7.
For each $p\geq 2$, given an orientation $\vec{G}$ of a $2p$-regular graph $G$, it is NP-complete to check whether there exists a $(p+2)$-star coloring $h$ of $G$ such that $(\vec{G},h)$ is a $(p+2)$-colored in-orientation. ∎
For $p\geq 3$, the complexity of checking whether a $2p$-regular graph is $(p+2)$-star colorable is posed as an open problem in [18]. Despite Theorem 3.7, this problem remains open.
By Theorems 3.3 and 3.2 (using $H=K_{q}$), we have the following.
Theorem 3.8.
Let $G$ be a connected graph of minimum degree at least 3. For each integer $q\geq 4$, there is a one-to-one correspondence between the set of OIHs from $\vec{L}(G)$ to $\vec{L}(K_{q})$ and set of $q$-colored MINI-orientations of $G$. ∎
We need a lemma for the next theorem. An orientation $\vec{H}$ is Eulerian if $\operatorname{out-degree}(v)=\operatorname{in-degree}(v)$ for every vertex $v$ of $\vec{H}$.
Lemma 3.
Let $G$ and $H$ be two $2p$-regular graphs, where $p\geq 2$. Let $\vec{G}$ be an orientation of $G$, and let $\vec{H}$ be an Eulerian orientation of $H$. Then, every OIH from $\vec{G}$ to $\vec{H}$ is an OBH from $\vec{G}$ to $\vec{H}$.
Proof.
Suppose that $\psi$ is an OIH from $\vec{G}$ to $\vec{H}$. Then, for every vertex $v$ of $\vec{G}$, $\operatorname{out-degree}(v)\leq p$ and thus $\operatorname{in-degree}(v)\geq p$; that is, $\operatorname{out-degree}(v)\leq p\leq\operatorname{in-degree}(v)$ for every vertex $v$. Since the sum of out-degrees equals the sum of in-degrees, $\operatorname{out-degree}(v)=p=\operatorname{in-degree}(v)$ for each vertex $v$ of $\vec{G}$. That is, $\vec{G}$ is Eulerian. Thus, for each vertex $v$ of $\vec{G}$, the sets $N^{+}_{\vec{G}}(v)$ and $N^{+}_{\vec{H}}(\psi(v))$ have the same cardinality. For each vertex $v$ of $\vec{G}$, since the restriction of $\psi$ to $N^{+}_{\vec{G}}(v)$ is an injection, the restriction must be a bijection. This shows that $\psi$ is an OBH from $\vec{G}$ to $\vec{H}$. ∎
An orientation $\vec{G}$ is a $2p$-regular graph admits a $(p+2)$-colored MINI-orientation if and only if $\vec{G}$ admits an OBH to $\vec{L}(K_{p+2})$ [3, Lemma 23] (a direct proof is presented in [2, Theorem 5.7], but with a different terminology; see para. 1 and Note 1 of [3, Page 18]). The term ‘OBH’ in the previous statement can be replaced by ‘OIH’, thanks to Lemma 3. Thus, we have the following.
Theorem 3.9.
For $p\geq 2$, given a $(p+1)$-regular graph $G$, it is NP-complete to check whether $\vec{L}(G)$ admits an OBH ((resp. OIH)) to $\vec{L}(K_{p+2})$. ∎
Theorem 3.10.
For $q\geq 4$, it is NP-complete to check whether a given orientation of a 2(q1)2(q-1)-regular graph admits an OBH ((resp. OIH)) to $\vec{L}(K_{q})$. ∎
4 Conclusion
In this work, we determined the characteristic polynomials of the adjacency matrix of $L^{*}(G)$ and the skew-symmetric adjacency matrix (and Hermitian adjacency matrix) of $\vec{L}(G)$ for $d$-regular graphs $G$ with $d\geq 3$, and presented a consequence to star coloring of regular graphs. In addition, we proved that there exists an LBH (resp. LIH) from a graph $G$ of minimum degree at least 3 to a graph $H$ if and only if there exists an OBH (resp. OIH) from $\vec{L}(G)$ to $\vec{L}(H)$. Thereby, we employed the oriented line graph operation to prove the first NP-completeness results on OBH and OIH. In general, the oriented line graph operation could be useful to link properties of undirected graphs to their directed versions (e.g. connectivity of graphs to connectivity of digraphs), especially when the focus is on the complexity of the properties.
Our result characterizing distance-two coloring in terms of colored MINI-orientations could be of interest from a different direction, namely understanding the relations between distance-two coloring and star coloring.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.